FEL Upgrade Project Weekly Brief
December 11-15, 2000


A draft version of the accelerator driver design (Version 1.1 ) was completed this week.


The project cost-performance report for the Month of November was completed this week.  The project remains essentially on cost and schedule with a positive schedule variance of $109k and a negative cost variance of $9k. 22%
of the project was scheduled to be complete by Nov.30 (assuming an April 1,2000 start), and we have completed 23% by Nov. 30.

WBS 3 (Beam Physics):

A proposal by W. Colson, NPGS, for accelerator physics and FEL studies was evaluated and will be funded.

A draft Revision 1.1 IR Upgrade machine design was completed and is undergoing performance evaluation studies. This revision uses "best knowledge" values for initial RF gradients (final energy of 145 MeV suspected) and consequently employs an enlarged injection condition (bx,y~10 m, ax,y=0) at 10 MeV (to compensate reduced focussing along the
linac) and asymmetric reinjection conditions at full energy (for similar reasons). The machine is slightly wider and shorter and uses slightly different main dipole angles than the Rev 1.0 design. This allows both additional space for correction packages in the end loops and allows future integration of additional FEL transport lines (such as the UV) through use of preexisting dipole designs.

WBS 4 (Injector):

Conditioned the titanium electrode in the Field Emission Test system up to 20.83 MV/m. Emission at 20 MV/m  on a 6mm gap was 84 nA. The titanium has preformed better than stainless steel but still not nearly as well as the nitrogen implanted stainless steel.

WBS 5 (SRF):

A current status report on the cavities for the new cryomodule:

The second and third production cavities are being completed (EBW) today.

The first cavity is currently being tuned. Cavity 4,5 cells are pressed and are waiting to be trimmed before EB welding. All flanges are machined.

All Beam tubes have had the nipples pulled for HOM tubes and are in chemistry for cleaning.  A copper seven cell has been setup to start HOM measurements and two HOM (copper) prototypes have been completed.

The first two Helium vessels have been received from the vendor and are under going QA.

WBS 6 (RF):

Zone 4 - No progress.  Testing is held up until the LCW is repaired.  A leak was discovered in the LCW transport line over to the FEL Bldg at the end of last week. The leak was repaired but the LCW loop within the FEL Bldg has to be cleaned up prior to any use of the LCW.

Zone 3 - No progress.  Waiting for additional parts.

WBS 9 (Transport):

• In anticipation of DOE’s approval of the Engineering Services contract, we composed the interface document and collected its attachments so that DULY Engineering could productively start the design of the Injection and Extraction dipoles (GU & GV).  This document will serve as a model for giving the design of the Arc Bend (UV Modified GX)
the Reverse Bend (GQ) and the 180 degree bend (GY) to the other contractor.
• During generation of the interface document above, it became evident that the requirements of the original IR Demo should be simplified for the upgrade.  These magnets could be cheaper.  New assumptions are:
• These magnets would not be used to divert partially accelerated, straight-ahead beam to the extraction dump for phasing.
• The margin of 20% over field could be decreased to 10% (11 MeV)
• With the above assumptions, solid magnet wire with plate cooling was substituted in the GU and GV for the hollow copper conductor used in the IR Demo.
• We received additional budgetary estimates for our dipoles and received cost saving construction method suggestions.
• Additional GW dipole detailing continued, concentrating on simplifying the bolting and alignment pin features as well as making the hoses larger and altering the magnet lead area in response to the thinner yoke material.
• We ordered the copper for the GW Prototype.

• QX (3.125” Quad) Magnets
 Prototype Fabrication and measurement: magnet assembly was completed and measurements started.  Reproducibility of the pole tip field was measured with a hall probe and repeated to +/-0.03%.  The gradient integral was measured with the existing 2" diameter rotating coil and found to meet the design spec.  Measurements will be made over the next few days on the 2" harmonic probe to get a feel for the field quality at the smaller radius.  Final field quality measurements, and the subsequent development of end chamfer modifications, will await the new 3" probe.
• Budgetary Quote: Several conversations were held with Magnet Enterprises International (MEI) and hopefully they will present a quote shortly.  A lack of response from two other "new" coil vendors is taken as a lack of interest and will not be pursued.
• Measurement Probe: Fabrication of the 100-turn litz coil is progressing.  Drawings for the probe body have been signed off and a local shop has been given the job.

• The Engineering Services Contracts remain in the DOE Office for approval.

WBS 10 (Wiggler):

PECO has started the rough machining of the pole pieces for the dispersion section.  They are expecting the first coil from their subcontractor the beginning of January.  Attachment of the modified pole clamps has started on the second half of the wiggler.  Design of the dispersion dipole mounting brackets has started. Next will be the design of the wiggler cover.  Finished the design of the wiggler viewer optical transport assembly.  All the major design work is done for the wiggler vacuum chamber and diagnostics.  Detailing should be done soon.  This leaves the dispersion section supports, the wiggler cover and the measurement jigs as the only things left to design.

WBS 11 (Optics):

Work at AES on the finite element analyses for the deformable mirror is on schedule. The optics group had an internal review of the conceptual design of the optical cavity components that was quite useful, these comments are being written into the specifications documentation and being passed on to the ME group as input for our designers.  We also had
a discussion on upgrades to the O-BPM system that will make them operate at wavelengths longer than 7 microns.  Two concepts were identified as being promising, and we are assembling hardware to try them.

Discussions were held with the two major vendors of laser interferometers (a diagnostic needed for the mirror test stand).  On the basis of these discussions we will begin the procurement of an instrument.


No operations were scheduled this week. Maintenance activities were done on various optical systems and the LCW system.

Privacy and Security Notice